# positioning

### v1

@@ 1.3. Holomorphic and VHS results
To enter in the world of variations of Hodge structures, we focus on complex hyperbolic lattices $\Gamma\subset G=\operatorname{\operatorname{PU}}(1,n)$ with $n>1$. See Section 5 for the various notions from Hodge theory used in this subsection. In this setting we can restate 1.7 in a more geometric way. Suppose for example that $H$ is the automorphism group of a Hermitian symmetric domain $X_{H}$, $\rho:\Gamma\to H$ takes values in a lattice $\Gamma_{H}\subset H$, and is induced by taking the induced map on fundamental groups for a holomorphic map
(1.3.1) $f:S_{\Gamma}\longrightarrow\Gamma_{H}\backslash X_{H}.$
@@
We remark here that the setting described in (1.3.1) is a special case of the variation of Hodge structure (VHS, from now on) formalism. The next theorem shows that rich representations underlying an integral pure polarized VHS extend.
@@
The proof of 1.14 builds on the main results of [7], namely the geometric part of the Zilber–Pink conjecture. In particular, the proof crucially uses the following, which we prove in Section 6, and may be of independent interest.
@@
1.4. Closed geodesics
To conclude the introduction, we study the one-dimensional case avoided in our previous results. From a dynamical point of view, the dimension at least two hypothesis is natural and critical in that it allows one to access tools from unipotent dynamics. From a Hodge theoretic point of view, one typically works with a complex subvariety. However, there is still something one can say for closed geodesics.
We begin with a simple cautionary example regarding generalizing richness to closed geodesics. Let $G$ be $\operatorname{\operatorname{PO}}(1,n)$ for $n\geq 3$ or $\operatorname{\operatorname{PU}}(1,n)$ for $n\geq 2$, and suppose that $\Gamma_{1},\Gamma_{2}$ are arithmetic lattices in $G$ of simplest type for which there is a surjection $\rho:\Gamma_{1}\to\Gamma_{2}$ with infinite kernel; such representations are known to exist in both settings. Choose a maximal subgroup $\Delta\subset\Gamma_{2}$ associated with a proper totally geodesic submanifold of $S_{\Gamma_{2}}$. Then $\rho^{-1}(\Delta)$ contains the kernel of $\rho$, and therefore is Zariski dense in $G$. Consequently, $\rho^{-1}(\Delta)$ contains a profinitely open collection of regular semisimple elements [34], hence they preserve no totally geodesic subspace of $X$ except their axis of translation, and thus the associated closed geodesics on $S_{\Gamma}$ are maximal. In the complex hyperbolic setting the associated closed geodesics are generic in the sense that their union is Zariski dense in $S_{\Gamma}$, hence this provides a fairly rich collection of $1$-dimensional maximal totally geodesic subspaces.
Representations as in the previous paragraph are unbounded and Zariski dense with connected, adjoint target, but they cannot extend since they have infinite kernel. They have a well-distributed sequence of maximal closed geodesics for which the associated elements in $\rho(\Gamma_{1})$ all preserve a totally geodesic subspace of the target, hence the representation could be considered rich in analogy with the higher-dimensional setting. While one might reasonably interpret this as saying the $1$-dimensional case is hopeless, our final result indicates that stronger geometric hypotheses can still lead to a rigidity theorem.
Theorem 1.16.
$f:S\longrightarrow\Gamma_{m}\backslash\mathbb{B}^{m}.$
Let $\{C_{\ell}\}$ be a generic sequence of closed geodesics in $S$. Assume that for each $\ell$ there exits a strict complex totally geodesic subvariety $Y_{\ell}\subset\Gamma_{m}\backslash\mathbb{B}^{m}$ such that $f(C_{\ell})\subset Y_{\ell}$. Then $f$ is a totally geodesic immersion.
See 4.4 for the proof. Note that 1.16 applies for nonarithmetic $\Gamma$, since all finite volume ball quotients contain infinitely many free homotopy classes of closed geodesics. That $f$ must take certain closed geodesics into a proper totally geodesic subspace is a very strong hypothesis, but it is not entirely unprecedented. Indeed one can show that some of the maps associated with surjections between complex hyperbolic lattices take a certain (necessarily finite) collection of totally geodesic subspaces of $S_{\Gamma_{n}}$ to totally geodesic subspaces of $S_{\Gamma_{m}}$. We close with a final question that would generalize 1.16.
Question 1.17.
1.5. Outline of paper
In Section 2 we prove the statement about compatibility. The rest of the paper is devoted to the complex hyperbolic case. In Section 3 we give the argument based on various forms of Ratner’s equidistribution theorem. In Section 4 we give a simple self-contained proof of the results about Siu’s immersion problem. The rest of the paper uses more ideas and techniques from Hodge theory. In particular Section 5 is devoted to all the needed preliminaries from Hodge theory and the Zilber–Pink paradigm. In Section 6 we prove the most general statement on isogenies of VHSs, namely 1.15. Finally we prove the main results about the extension of rich representations underlying a VHS, see Section 7.
@@ 2. Dynamics background and compatibility
This section contains some background on the dynamical side, along with the proof of the first part of 1.11. Let $G$ be either $\operatorname{\operatorname{PO}}(1,n)$ or $\operatorname{\operatorname{PU}}(1,m)$. We first state some basic facts and establish some notation that will be used throughout this paper. What is described immediately below recaps facts that are given a detailed treatment in [1, 2].
Fixing a maximal compact subgroup $K$ in $G$, the associated locally symmetric space is $X=G/K$. Fix a standard representative $Y\subset X$ for each type (i.e., $G$-orbit) of totally geodesic subspace of $X$ with dimension at least two. If $W$ is the subgroup of the stabilizer of $Y$ in $G$ generated by unipotent elements, then the full stabilizer of $Y$ is the normalizer $N_{G}(W)$, which is a compact extension of $W$.
@@
Since it arises several times, we briefly recall what it means for $\Gamma\subset G$ to be an arithmetic lattice of the simplest kind. These are well-studied in the literature, so we are terse. In the real hyperbolic case, this means that there is a totally real number field $F$ with integer ring $\mathcal{O}_{F}$ and a quadratic form $q$ on $F^{n+1}$ so that $\Gamma$ is commensurable with $\mathcal{G}(\mathcal{O}_{F})$, where $\mathcal{G}$ is the $F$-algebraic group $\operatorname{\operatorname{PO}}(q)$. In the complex hyperbolic case, there is a totally imaginary extension $E$ of the totally real field $F$ so that $\Gamma$ is commensurable with $\mathcal{G}(\mathcal{O}_{F})$, where $\mathcal{G}$ is now $\operatorname{\operatorname{PU}}(h)$ for a hermitian form $h$ on $E^{n+1}$ (which is still an $F$-algebraic group). In each case, it is necessarily the case that the form has signature $(1,n)$ at one place of $F$ and is definite at all other places.
@@
Suppose that $\Gamma\subset G$ is an arithmetic lattice so that the locally symmetric space $S_{\Gamma}$ contains a proper immersed totally geodesic subspace of dimension at least two. Let $k$ be a local field of characteristic zero and ${\mathbf{H}}$ be a connected, adjoint $k$-algebraic group so that the pair $(k,{\mathbf{H}})$ is compatible with $G$. If $\rho:\Gamma\to{\mathbf{H}}(k)$ is a rich representation, then $\rho$ extends.
@@
By 1.6, $\rho$ is a homomorphism with unbounded, Zariski dense image. Following Theorem 1.6 and Remark 1.7 of [1], to prove that there exists a continuous extension ${\widetilde{\rho}:G\to{\mathbf{H}}(k)}$ of $\rho$, it suffices to find
@@
    a $k$-rational faithful, irreducible representation ${\mathbf{H}}(k)\to\mathbf{SL}(V)$ on a finite dimensional $k$-vector space $V$, and
@@
    a $W$-invariant measure $\nu$ on $\Gamma\backslash(G\times\mathbb{P}(V))$ that projects to Haar measure $\mu_{\Gamma\backslash G}$ on $\Gamma\backslash G$.
@@
Recall that by hypothesis, $S_{\Gamma}$ contains infinitely many maximal totally geodesic subspaces $S_{\Gamma_{i}}$. Passing to a subsequence, we assume they are of the same type (in the sense described in this section). This defines a sequence of homogeneous, $W$-ergodic measures $\{\mu_{i}\}$ with $\mathrm{supp}(\mu_{i})\neq\Gamma\backslash G$ for which $\mu_{i}\to\mu_{\Gamma\backslash G}$, where $W\subset G$ is a fixed proper subgroup generated by unipotent elements. This provides the $W$ in item (1) above.
@@
Returning to the proof at hand, recall from 1.6 that ${\mathbf{H}}_{i}=\overline{\rho(\Gamma_{i})}$ is a $k$-defined noncompact proper subgroup of ${\mathbf{H}}$ with $\dim({\mathbf{H}}_{i})<\dim({\mathbf{H}})$ for infinitely many $i$. Passing to a subsequence, we can assume that $\dim({\mathbf{H}}_{i})$ is constant, and we denote this constant by $d$. Consider the $d^{th}$ exterior power of the adjoint representation
@@
where $\mathfrak{h}$ is the Lie algebra of ${\mathbf{H}}(k)$. The Lie algebra $\mathfrak{h}_{i}=\operatorname{Lie}({\mathbf{H}}_{i}(k))$ determines a line $\ell_{i}\subset\bigwedge\nolimits^{d}\mathfrak{h}$. The representation $\wedge^{d}(\operatorname{Ad})$ need not be irreducible, however there exists a nontrivial summand $\theta:{\mathbf{H}}(k)\to\mathbf{GL}(V)$, where $V$ is a $k$-defined vector space, for which infinitely many $\ell_{i}$ project nontrivially to $V$. This representation satisfies item (2) above.
@@
to another arithmetic ball quotient of simplest kind. Suppose that $Y_{\ell}$ is a sequence of irreducible totally geodesic divisors on $S_{m}$ for which the top dimensional irreducible components of $f^{-1}(Y_{\ell})$ are totally geodesic subvarieties of $S_{n}$. Then $f$ is a totally geodesic immersion.
@@
A simple transversality argument implies that $f^{-1}(Y_{\ell})$ has top-dimensional irreducible components with codimension one for all but finitely many $\ell$. Let $\beta$ be a positive $(n-1,n-1)$ form on $S_{m}$ with compact support. The $Y_{\ell}$ equidistribute in $S_{m}$, so by applying [27, Cor. 1.5] [^4] Also see the related paper of Möller–Toledo’s [30] and the more general work of Tayou–Tholozan [39]. to both $\beta$ and $\omega^{n-1}$ we deduce that
(3.0.1) $\frac{\int_{f(S_{n})\cap Y_{\ell}}\beta}{\int_{f(S_{n})\cap Y_{\ell}}\omega^{n-1}}\longrightarrow\frac{\int_{f(S_{n})}\beta\wedge\omega}{\int_{f(S_{n})}\omega^{n}},$
as $\ell$ goes to infinity. By definition of the pullback, the left hand side of Equation (3.0.1) is also equal to
@@
We want to compute the limit of $\frac{\int_{f^{-1}(Y_{\ell})}f^{*}\beta}{\int_{f^{-1}(Y_{\ell})}\alpha^{n-1}}$ and of $\frac{\int_{f^{-1}(Y_{\ell})}f^{*}\omega^{n-1}}{\int_{f^{-1}(Y_{\ell})}\alpha^{n-1}}$.
@@
Let $\{\mu_{f^{-1}(Y_{\ell})}\}$ be the following sequence of probability measures on $\Gamma_{n}\backslash G$
$\mu_{f^{-1}(Y_{\ell})}\coloneqq\frac{1}{n_{\ell}}\sum_{i=1}^{n_{\ell}}\mu_{Z_{\ell,i}}.$
Then the $\mu_{f^{-1}(Y_{\ell})}$ weak-normal-$\ast$ converge to $\mu_{S_{n}}$.
@@
We need to recall the description of the totally geodesic divisors on $S_{n}=\Gamma_{n}\backslash\mathbb{B}^{n}$. We refer to [2, Lem. 8.2] for a more detailed description and for a proof of the following facts. We have an isomorphism $\mathbb{B}^{n}=G/K$, where $K=\mathrm{S}(\operatorname{\operatorname{U}}(1)\times\operatorname{\operatorname{U}}(n))$ is a maximal compact subgroup of $G$, and a fixed embedding $W\coloneqq\operatorname{\operatorname{SU}}(1,n-1)\subset G$. Let $\pi:\Gamma_{n}\backslash G\longrightarrow S_{n}$ be the natural projection associated to the right quotient by $K$. Then any totally geodesic divisor $Z$ on $S_{n}$ is of the form
@@
Note that $\pi_{*}\mu_{\ell}=\mu_{f^{-1}(Y_{\ell})}$ by [1, Lem. 8.2(1)]. As each $\mu_{\ell,i}$ is $W$-invariant so is $\mu_{\ell}$ and consequently any weak-$\ast$ limit $\mu$ of a subsequence of the $\mu_{\ell}$ is a $W$-invariant measure. Moreover, the similar argument to [1, Lem. 8.3] using quantitative non-divergence shows that there is no escape of mass and $\mu$ is also a probability measure.
Let $\nu$ be an ergodic component of $\mu$. By Ratner’s theorem [35], $\nu$ is a homogeneous, $W$-invariant, $W$-ergodic measure and therefore is either a constant multiple of Haar measure on $\Gamma_{n}\backslash G$ or a constant multiple of Haar measure on a proper closed $W$-orbit $V\subsetneq\Gamma_{n}\backslash G$. However, note that for any such closed $W$-orbit, $\mu_{\ell,i}(V)=0$ for sufficiently large $\ell$ and all $1\leq i\leq n_{\ell}$ and hence the latter cannot occur. Therefore only the former possibility occurs and hence $\nu=\mu_{\Gamma_{n}\backslash G}$. Since pushforward and weak-$\ast$ limits commute, we conclude that $\mu_{f^{-1}(Y_{\ell})}\to\mu_{S_{n}}$ as required. ∎
As a consequence, again by upgrading Ratner to a statement about forms as in [27, Cor. 1.5], but this time in $S_{n}$ rather than in $\mathbb{B}^{m}$, we get
@@
is an immersive holomorphic map to another arithmetic ball quotient. Let $\Gamma_{f}$ be the graph of $f$ in $S_{n}\times S_{m}$. We now prove the following equivalent formulation of 1.3.
@@
Let $\{Z_{\ell}\}$ (resp. $\{Y_{\ell}\}$) be an infinite sequence of complex totally geodesic subvarieties of $S_{n}$ (resp. $S_{m}$). If $f(Z_{\ell})\subset Y_{\ell}$ for all $\ell$, then $f$ is a totally geodesic immersion.
@@
We can and do assume that $f(S_{n})$ is Hodge generic, i.e., is not contained in any strict complex totally geodesic subvariety $S^{\prime}$ of $S_{m}$, as $S^{\prime}$ would be a ball quotient of dimension $m^{\prime}<m$ and the result for $S_{m}$ is implied immediately if we prove it instead for $S^{\prime}$.
@@
(4.0.1) $\operatorname{codim}_{S_{n}\times S_{m}}(U_{\ell})<\operatorname{codim}_{S_{n}\times S_{m}}(\Gamma_{f})+\operatorname{codim}_{S_{n}\times S_{m}}(Z_{\ell}\times Y_{\ell}).$
@@
We can pass to a subsequence and assume that the dimensions of $Z_{\ell}$ and $Y_{\ell}$ are independent of $\ell$. Set $d=\dim Z_{\ell}$ and $\dim Y_{\ell}=m-e$. We then have
@@
Since $e>0$, Equation (4.0.1) holds. ∎
Applying 4.2 to $\Gamma_{f}\subset S_{n}\times S_{m}$, we conclude that $\Gamma_{f}$ is contained in a strict totally geodesic subvariety $S_{0}$ of $S_{n}\times S_{m}$. As $f(S_{n})$ is Hodge generic, the second projection of $S_{0}$ on $S_{m}$ is surjective. As $\Gamma_{f}$ is isomorphic to $S_{n}$ under the first projection, we deduce that $S_{0}$ is also isomorphic to $S_{n}$ by the first projection and that $f$ is a totally geodesic embedding. ∎
Using a similar argument we can also prove the following.
@@
to another finite-volume ball quotient. Let $\{C_{\ell}\}$ be a generic sequence of closed geodesics in $S_{n}$ for which there exits a strict complex totally geodesic subvariety $Y_{\ell}\subset S_{m}$ such that $f(C_{\ell})\subset Y_{\ell}$ for all $\ell$. Then $f$ is a totally geodesic immersion.
The proof of 4.4 uses the so called Baby Ax–Lindemann theorem [41, Prop. 2.6], which we recall below in the special case we need.
@@
Note that $S_{m}$ must be arithmetic. Indeed, if $S_{m}$ were nonarithmetic the fact that $\{C_{\ell}\}$ is generic implies that, after passing to maximal proper totally geodesic subvarieties, $f(S_{n})$ is contained in a finite union of totally geodesic subvarieties. Since the map is a smooth immersion, the image is contained in a single such subvariety $S^{\prime}$. Replacing $S_{m}$ with $S^{\prime}$, we see that $S^{\prime}$ must have infinitely many distinct maximal totally geodesic subvarieties, and hence is arithmetic.

### v3

@@ 1.3. Holomorphic and VHS results
To enter in the world of variations of Hodge structures, we focus on complex hyperbolic lattices $\Gamma\subset G=\operatorname{\operatorname{PU}}(1,n)$ with $n>1$. See Section 5 for the various notions from Hodge theory used in this subsection. In this setting we can restate 1.9 in a more geometric way. Suppose for example that $H$ is the automorphism group of a Hermitian symmetric domain $X_{H}$, $\rho:\Gamma\to H$ takes values in a lattice $\Gamma_{H}\subset H$, and is induced by taking the induced map on fundamental groups for a holomorphic map
(1.2) $f:S_{\Gamma}\longrightarrow\Gamma_{H}\backslash X_{H}.$
@@
We remark here that the setting described in (1.2) is a special case of the variation of Hodge structure (VHS, from now on) formalism. The next theorem shows that rich representations underlying an integral pure polarized VHS extend.
@@
The proof of 1.17 builds on the main results of [7], namely the geometric part of the Zilber–Pink conjecture. In particular, the proof crucially uses the following, result proved in Section 6, which may be of independent interest.
@@
1.4. Outline of the paper
In Section 2 we prove the statement about compatibility. The rest of the paper is devoted to the complex hyperbolic case. In Section 3 we give the argument based on various forms of Ratner’s equidistribution theorem. In Section 4 we give a simple self-contained proof of the results about Siu’s immersion problem. The rest of the paper uses more ideas and techniques from Hodge theory. In particular Section 5 is devoted to all the required preliminaries from Hodge theory and the Zilber–Pink paradigm. In Section 6 we prove the most general statement on isogenies of VHSs, namely 1.18. Finally we prove the main results about the extension of rich representations underlying a VHS in Section 7.
@@ 2. Dynamics background and compatibility
This section contains some background on the dynamical side, along with the proof of the first part of 1.13. Let $G$ be either $\operatorname{\operatorname{PO}}(1,n)$ or $\operatorname{\operatorname{PU}}(1,m)$. We first state some basic facts and establish some notation that will be used throughout this paper. What is described immediately below recaps facts that are given a detailed treatment in [1, 2].
Fixing a maximal compact subgroup $K$ in $G$, the associated locally symmetric space is $X=G/K$. Fix a representative $Y\subset X$ for each type (i.e., $G$-orbit) of totally geodesic subspace of $X$ with dimension at least two and call such a representative a standard subspace. If $W$ is the subgroup of the stabilizer of $Y$ in $G$ generated by unipotent elements, then the full stabilizer of $Y$ is the normalizer $N_{G}(W)$, which is a compact extension of $W$.
@@
Suppose that $\Gamma\subset G$ is a geodesically rich lattice. Let $k$ be a local field of characteristic zero and ${\mathbf{H}}$ be a connected, adjoint $k$-algebraic group so that the pair $(k,{\mathbf{H}})$ is compatible with $G$. If $\rho:\Gamma\to{\mathbf{H}}(k)$ is a rich representation, then $\rho$ extends.
@@
By 1.8, $\rho$ is a homomorphism with unbounded, Zariski dense image. Following Theorem 1.6 and Remark 1.7 of [1], to prove that there exists a continuous extension ${\widetilde{\rho}:G\to{\mathbf{H}}(k)}$ of $\rho$, it suffices to find
@@
    a $k$-rational faithful, irreducible representation ${\mathbf{H}}\to\mathbf{SL}(V)$ on a finite dimensional $k$-vector space $V$, and
@@
    a $W$-invariant measure $\nu$ on $\Gamma\backslash(G\times\mathbb{P}(V))$ that projects to the Haar measure $\mu_{\Gamma\backslash G}$ on $\Gamma\backslash G$.
@@
Recall that by hypothesis, $S_{\Gamma}$ contains infinitely many distinct maximal totally geodesic subspaces $S_{\Gamma_{i}}$. Passing to a subsequence, we assume they are of the same type (in the sense described in this section). This defines a sequence of homogeneous, $W$-ergodic measures $\{\mu_{i}\}$ with $\mathrm{supp}(\mu_{i})\neq\Gamma\backslash G$ for which $\mu_{i}\to\mu_{\Gamma\backslash G}$, where $W\subset G$ is a fixed proper subgroup generated by unipotent elements. This provides the $W$ in item (1) above.
@@
Returning to the proof at hand, recall from 1.8 that ${\mathbf{H}}_{i}=\overline{\rho(\Gamma_{i})}$ is a $k$-defined noncompact proper subgroup of ${\mathbf{H}}$ with $\dim({\mathbf{H}}_{i})<\dim({\mathbf{H}})$ for infinitely many $i$. Passing to a subsequence, we can assume that $\dim({\mathbf{H}}_{i})$ is constant, and we denote this constant by $d$. Consider the $d^{th}$ exterior power of the adjoint representation
@@
where $\mathfrak{h}$ is the Lie algebra of ${\mathbf{H}}(k)$. The Lie algebra $\mathfrak{h}_{i}=\operatorname{Lie}({\mathbf{H}}_{i}(k))$ determines a line $\ell_{i}\subset\bigwedge\nolimits^{d}\mathfrak{h}$. The representation $\wedge^{d}(\operatorname{Ad})$ need not be irreducible, however there exists a nontrivial summand $\theta:{\mathbf{H}}\to\mathbf{GL}(V)$, where $V$ is a $k$-defined vector space, for which infinitely many $\ell_{i}$ project nontrivially to $V$. This representation satisfies item (2) above.
@@
to another arithmetic ball quotient of simplest kind. Suppose that $Y_{\ell}$ is a sequence of distinct irreducible totally geodesic divisors on $S_{m}$ for which the top dimensional irreducible components of $f^{-1}(Y_{\ell})$ are totally geodesic subvarieties of $S_{n}$. Then $f$ is a totally geodesic immersion.
@@
A simple transversality argument implies that $f^{-1}(Y_{\ell})$ has top-dimensional irreducible components with codimension one for all but finitely many $\ell$. Let $\beta$ be a positive $(n-1,n-1)$ form on $S_{m}$ with compact support. The $Y_{\ell}$ equidistribute in $S_{m}$, so by applying [30, Cor. 1.5] [^4] Also see the related paper of Möller–Toledo’s [33] and the more general work of Tayou–Tholozan [42]. to both $\beta$ and $\omega^{n-1}$ we deduce that
(3.1) $\frac{\int_{f(S_{n})\cap Y_{\ell}}\beta}{\int_{f(S_{n})\cap Y_{\ell}}\omega^{n-1}}\longrightarrow\frac{\int_{f(S_{n})}\beta\wedge\omega}{\int_{f(S_{n})}\omega^{n}}$
as $\ell$ goes to infinity. By definition of the pullback, the left hand side of Equation (3.1) is also equal to
@@
We want to compute the limits of $\frac{\int_{f^{-1}(Y_{\ell})}f^{*}\beta}{\int_{f^{-1}(Y_{\ell})}\alpha^{n-1}}$ and $\frac{\int_{f^{-1}(Y_{\ell})}f^{*}\omega^{n-1}}{\int_{f^{-1}(Y_{\ell})}\alpha^{n-1}}$.
@@
Let $\{\mu_{f^{-1}(Y_{\ell})}\}$ be the sequence
$\mu_{f^{-1}(Y_{\ell})}\coloneqq\frac{1}{n_{\ell}}\sum_{i=1}^{n_{\ell}}\mu_{Z_{\ell,i}}$
of probability measures on $\Gamma_{n}\backslash G$. Then the $\mu_{f^{-1}(Y_{\ell})}$ weak-$\ast$ converge to $\mu_{S_{n}}$.
@@
We need to recall the description of the totally geodesic divisors on $S_{n}=\Gamma_{n}\backslash\mathbb{B}^{n}$. We refer to [2, Lem. 8.2] for a more detailed description and for a proof of the following facts. We have an isomorphism $\mathbb{B}^{n}=G/K$, where $K=\mathrm{S}(\operatorname{\operatorname{U}}(1)\times\operatorname{\operatorname{U}}(n))$ is a maximal compact subgroup of $G$, and a fixed embedding $W\coloneqq\operatorname{\operatorname{SU}}(1,n-1)\subset G$. Let $\pi:\Gamma_{n}\backslash G\longrightarrow S_{n}$ be the natural projection associated with the right quotient by $K$. Then any totally geodesic divisor $Z$ on $S_{n}$ is of the form
@@
Note that $\pi_{*}\mu_{\ell}=\mu_{f^{-1}(Y_{\ell})}$ by [1, Lem. 8.2(1)]. As each $\mu_{\ell,i}$ is $W$-invariant so is $\mu_{\ell}$ and consequently any weak-$\ast$ limit $\mu$ of a subsequence of the $\mu_{\ell}$ is a $W$-invariant measure. Moreover, the similar argument to [1, Lem. 8.3] using quantitative non-divergence shows that there is no escape of mass and thus $\mu$ is also a probability measure.
Let $\nu$ be an ergodic component of $\mu$. By Ratner’s theorem [38], $\nu$ is a homogeneous, $W$-invariant, $W$-ergodic measure and therefore is either a constant multiple of the Haar measure on $\Gamma_{n}\backslash G$ or a constant multiple of the Haar measure on a proper closed $W$-orbit $V\subsetneq\Gamma_{n}\backslash G$. However, note that for any such closed $W$-orbit, $\mu_{\ell,i}(V)=0$ for sufficiently large $\ell$ and all $1\leq i\leq n_{\ell}$ and hence the latter cannot occur. Therefore only the former possibility occurs and hence $\nu=\mu_{\Gamma_{n}\backslash G}$. Since pushforward and weak-$\ast$ limits commute, we conclude that $\mu_{f^{-1}(Y_{\ell})}\to\mu_{S_{n}}$ as required. ∎
As a consequence, again by upgrading Ratner to a statement about forms as in [30, Cor. 1.5], but this time in $S_{n}$ rather than in $\mathbb{B}^{m}$, we get
@@
is an immersive holomorphic map to another arithmetic ball quotient. From now on, we always assume that $f(S_{n})$ is monodromy generic, i.e., is not contained in any strict complex totally geodesic subvariety $S^{\prime}$ of $S_{m}$. Let $\Gamma_{f}$ be the graph of $f$ in $S_{n}\times S_{m}$. We now prove the following equivalent formulation of 1.3.
@@
Let $\{Z_{\ell}\}$ (resp. $\{Y_{\ell}\}$) be an infinite sequence of distinct complex totally geodesic subvarieties of $S_{n}$ (resp. $S_{m}$). If $f(Z_{\ell})\subset Y_{\ell}$ for all $\ell$, then $f$ is a totally geodesic immersion.
@@
(4.1) $\operatorname{codim}_{S_{n}\times S_{m}}(U_{\ell})<\operatorname{codim}_{S_{n}\times S_{m}}(\Gamma_{f})+\operatorname{codim}_{S_{n}\times S_{m}}(Z_{\ell}\times Y_{\ell}).$
@@
We can pass to a subsequence and assume that the dimensions of $Z_{\ell}$ and $Y_{\ell}$ are independent of $\ell$. Set $d=\dim Z_{\ell}$ and $\dim Y_{\ell}=m-e$. We then have:
@@
Since $e>0$, Equation (4.1) holds. ∎
Since we assumed that $f(S_{n})$ is monodromy generic, we must prove that $f(S_{n})=S_{m}$. Applying 4.2 to $\Gamma_{f}\subset S_{n}\times S_{m}$, we conclude that $\Gamma_{f}$ is contained in a strict totally geodesic subvariety $S_{0}$ of $S_{n}\times S_{m}$. As $f(S_{n})$ is monodromy generic, the second projection of $S_{0}$ on $S_{m}$ is surjective. Therefore $S_{0}$ is a strict totally geodesic subvariety of $S_{n}\times S_{m}$ dominating both factors. This is possible only when $S_{n}$ is commensurable with $S_{m}$ and $S_{0}$ is a modular correspondence. Since $\Gamma_{f}\subset S_{0}$ have the same dimension and are irreducible, it follows that $\Gamma_{f}=S_{0}$, i.e., $f$ is a totally geodesic embedding. ∎
Using a similar argument we can also prove the following, where a sequence of geodesics is generic when they are not contained in a finite union of maximal totally geodesic subvarieties.
@@
to another finite-volume ball quotient. Let $\{C_{\ell}\}$ be a generic sequence of distinct closed geodesics in $S_{n}$ for which there exists a strict complex totally geodesic subvariety $Y_{\ell}\subset S_{m}$ such that $f(C_{\ell})\subset Y_{\ell}$ for all $\ell$. Then $f$ is a totally geodesic immersion.
The proof of 4.4 uses the so called Baby Ax–Lindemann theorem [44, Prop. 2.6], which we recall below in the special case we need.
@@
Note that we may reduce to the case that $S_{m}$ is arithmetic. Indeed, if $S_{m}$ were nonarithmetic the fact that $\{C_{\ell}\}$ is generic implies that, after passing to maximal proper totally geodesic subvarieties, $f(S_{n})$ is contained in a finite union of totally geodesic subvarieties. Since the map is a smooth immersion, the image is contained in a single such subvariety $S^{\prime}$. Replacing $S_{m}$ with $S^{\prime}$, we see that $S^{\prime}$ must have infinitely many distinct maximal totally geodesic subvarieties (see [2] or [8]), and hence is arithmetic.